re of length 0.35 m is conducting a current of I 8.7 A in the +x direction through a B 5.0-T uniform magnetic field that is directed parallel to the wire in the -x direction. What are the magnitude and direction of the magnetic force on the wire? Number +X O +y O -y none of the above BINEED BOTH PARTS
Answer #1

a] Force on a current carrying wire due to a magnetic field is given by: F = BILsin\theta

where \theta is the angle between the wire and the external magnetic field.

for the first part: \theta = 0o and so F = BILsin(0) = 0 N

b] \theta = 90o

6 = B(8.7)(0.35)sin90o

=> B = 1.97 T in the negative z direction.